# Risk-Calibrated Human-Robot Interaction via Set-Valued Intent Prediction

### v1

@@ IV-E Multi-Step, Multi-Risk Control
$\mathbb{P}^{(\tilde{x},\bar{z})\sim\bar{\mathcal{D}}}(\text{OOD})=\mathbb{P}^{(\tilde{x},\bar{z})\sim\bar{\mathcal{D}}}\Big{(}u^{R*}(z)\notin\mathcal{T}_{\phi}(\bar{{x}})\wedge|\mathcal{T}_{\phi}(\bar{{x}})|\leq 1\Big{)},$ (23)
@@
In the multi-step, multi-risk setting, we consider a set of sequence-level risk signals $(R_{1},...,R_{k})$ for contexts in $\bar{\mathcal{D}}$ bounded at nominal levels $(\alpha_{1},...,\alpha_{k})$ by all $\phi\in\hat{\Phi}$ as before. We assume that each risk models an event $E_{k}$, and the loss for each risk $L_{k}$ is an indicator function $\mathbbm{1}[E_{k}]$. We assume that $R_{1}=R_{\text{cov}}$. In addition, since any OOD sequence incurs a task failure, we seek to bound the probability of $E_{k}$ occurring subject to an $R_{1}$ probability of distribution shift (in which case $E_{k}$ can also happen).
@@
[…]
@@
$\alpha_{1}+\alpha_{k}$ $\geq R_{1}(\phi)+R_{k}(\phi)$ (24)
$=\mathbb{E}^{(\tilde{x},\bar{z})\sim\bar{\mathcal{D}}}\Big{[}L_{1}(\tilde{x},\mathcal{T}_{\phi}(\tilde{x}),\bar{z})+L_{k}(\tilde{x},\mathcal{T}_{\phi}(\tilde{x}),\bar{z})\Big{]}$
@@ V Experiments
Environments. We demonstrate RCIP in three multi-step, interactive domains, which exhibit three ways in which a robot planner can be integrated with an intent predictor. First, we consider a multi-hallway setting in which an autonomous vehicle and a human-driven vehicle must coordinate to reach opposite ends of a room by navigating a set of hallways that are only one vehicle-width wide (see Fig. 3). The human vehicle randomly selects one of the hallways in advance but does not communicate the hallway to the robot. Next, we investigate human-robot cooperative navigation in close-quarters, cluttered household settings in the Habitat 3.0 [44] simulator (see Fig. 4). Finally, we show simulation and hardware experiments for zero-shot cooperative manipulation, in which the robot aids the human in sorting common household objects (e.g. books, toys, and fruit) by a mixture of size, shape, and color (see Fig. 1). Since the environment dynamics (1) may evolve at a much faster time scale than the human’s intent dynamics (2), the human’s intent is updated once every $T_{z}$ timesteps and is constant otherwise.
@@
Scenario Distribution and Calibration Dataset. RCIP can be used to obtain risk guarantees for an unknown scenario distribution — that is, of environments and human partners — if can can collect i.i.d. samples from it for calibration. We envision that RCIP will enable a robot to interact with an end user (or set of users) through interactive data collection. Then, using the set of FWER-controlling parameters obtained from calibration (cf. section IV), the user may set a level of autonomy for the robot depending on the risk limits of the task. The scenario distribution for each environment is described in the following subsections. Each calibration dataset is generated by random sampling from the environment distribution and from the distribution over human intents. For the simulation environments, we use a pre-trained prediction model using $10k$ random scenarios. For the hallway and cooperative navigation environments, the prediction models are trained on a single NVIDIA GeForce RTX 2080 Ti GPU. Pre-training the prediction model takes about $4$ hours per environment. For calibration on hardware, data collection takes about $8$ hours. For all environments, we fix $\delta=0.01$ and use a calibration dataset of size $M=500$. In all experiments, we evaluate thresholds $\lambda\in[0,1]$ with a step size of $0.001$ and model temperature $\theta\in\{0.2,0.4,0.6,0.8,1\}$.
@@
Baselines. We compare RCIP against similar set-valued prediction approaches. A simple but naive approach for approximated $1-\alpha_{1}$ coverage of optimal actions is Simple Set, which ranks actions according to a $1-\alpha_{1}$ threshold using the predictor’s raw confidence scores. Actions are sorted by greatest to least confidence, and actions are added to the prediction set in order of the sorted action set until the threshold is reached. To measure the effect of overall uncertainty rather than individual scores, we compare against Entropy Set, which includes the highest overall prediction if the entropy of the distribution predicted actions is below a threshold; if not, then all actions are included in the prediction set, and the robot must ask for help. To evaluate the performance of vanilla conformal prediction against the richer hypothesis space of RCIP, we report results for KnowNo [36]. Similar in spirit but different from our work, KnowNo seeks to maximize coverage of optimal actions but without any guarantees on the human help rate, and assumes model parameters are fixed. Instead of maximizing coverage outright, RCIP balances prediction of optimal actions with limits on the human help rate, providing flexible performance guarantees depending on model parameters. Lastly, we consider No Help as an option, where the predicted action set always contains the predictor’s most-confident action, and the human help rate is identically zero.
Figure 4: Multi-step RCIP is applied in Social Navigation. The human and robot are tasked with finding and collecting a series of objects (e.g. cans of soup, crackers) around a close-quarters living space. The robot must recognize the human’s intent and either follow or evade the human depending on the human’s desired object. The robot must minimize action miscoverage across a variety of human intents and environments.
@@
V-A Simulation: Multi-Vehicle Autonomous Navigation
Figure 5: Baseline comparison for RCIP verus other set-valued predictors for Hallway Navigation. RCIP provides a framework for tuning model parameters to achieve risk control, versus other methods that assume that model parameters are held fixed: KnowNo [36], Simple Set, Entropy Set, and No Help.
Autonomous navigation around other autonomous decision-making agents, including humans, requires the robot to recognize scenario uncertainty (whether another agent will turn right or left) with task efficiency (energy spent braking or taking detours). While safety can almost always be guaranteed if each vehicle declares their intent at all times, such communication can be costly, especially if human prompting is involved. In this example (Fig. 3), the robot is asked to navigate to the human vehicle’s initial condition without colliding while the human does the same. The set of intents is $\mathcal{Z}=\{1,2,3,4,5\}$, where each intent corresponds to one of the five hallways. The confidence scores for each intent are computed by taking the temperature-weighted softmax scores for each hallway. The final action probabilities are computed according to Eqn. (10). The robot interacts with the human over $T=200$ environment time steps and predicts the human’s intent every $T_{z}=20$ time steps.
To ensure that the robot reaches its goal state in a minimal amount of time, we permit the robot to prompt the human for their chosen hallway if its optimal action set is not a singleton. We jointly learn the robot and human policies using proximal policy optimizatio (PPO) [45, 46]. The human and robot PPO policies are trained jointly using $256$ environments and take about $4$ hours to train.
@@ V-A Simulation: Multi-Vehicle Autonomous Navigation
RCIP $0.85$ $0.86$ 0.34 0.95 0.24
KnowNo [36] $0.85$ $0.86$ $0.48$ $0.92$ $0.42$
Simple Set $0.98$ $0.85$ $0.48$ $0.92$ $0.42$
Entropy Set -- $0.75$ $0.07$ $0.86$ $0.02$
No Help -- $0.73$ 00 $0.86$ 00
@@
TABLE I: Results for Cooperative Navigation. The optimal action miscoverage rate is held fixed between RCIP, KnowNo, and Simple Set for comparing the other metrics.
V-B Simulation: Cooperative Navigation in Habitat
@@
Habitat [44] is a photo-realistic simulator containing a diverse set of scenes, objects, and humans models for human robotics tasks. In this experiment, a Boston Dynamics Spot robot and human are jointly tasked with navigating to a set of goal objects in sequence, to simulate cleaning up a house (i.e., grabbing various items, such as crackers, cans of soup, etc. as shown in Fig. 4). Each scene contains $5-10$ objects of interest. Although the human may initially be out of view of the robot, the robot must find the human and maintain a safe distance of one meter at all times. We simulate the human’s decision making by choosing a high-level intent from the set of objects; here, $\mathcal{Z}=[N_{o}]$, where $N_{o}$ is the number of objects in the scene. The confidence scores for each intended object are computed by taking the temperature-weighted softmax scores for each goal object. The final action probabilities are computed according to Eqn. (10). The robot interacts with the human over $T=600$ environment time steps and selects a new goal object every $T_{z}=100$ time steps.
@@
Since the human’s goal object is not observed by the robot, one naive strategy is to navigate to the human first, then follow the human around the house. However, since the scene is cluttered, remaining too close to the human could impede their progress (e.g. getting in the way) or block the robot, resulting in suboptimal, unsafe behavior. By predicting the human’s motion, the robot is able to better accommodate the human’s task while remaining safe (with high probability) with respect to unsafe interactions. We present results for cooperative navigation in Table I.
V-C Hardware: Cooperative Manipulation
@@
In this example (Fig. 1), each scenario tasks the robot with helping a human to sort a set of objects by inferring the sorting category for each object. Since the human may have a preference for how the robot sorts the objects, the robot is additionally tasked with exactly matching the human’s sorting preferences (e.g. if the human wants the objects sorted by color, then the robot cannot add a conflicting color to the pile). We assume that the human’s intent set $\mathcal{Z}$ is represented in the (high-dimensional) space of natural language descriptions, such as “the color orange”, “children’s toys”, and “vegetables”, and that intents and actions are one-to-one. The robot interacts with the human over $T=8$ environment time steps, and the human selects a new sorting plan every $T_{z}=1$ time step.
@@
To simplify the planning task, we assume that the robot takes in an image observation of the table and has access to a (vision) language model to process the semantic features of the image. We first use GPT-4V (gpt-4-vision-preview) [3] to process the image by asking for a description of each bin and the item to be sorted (e.g. the carrot in Fig. 1), commenting on possible sorting criteria for each bin. Then, using the bin descriptions, we prompt a language-only model (gpt-3.5-turbo) to rank a set of possible plans via multiple-choice question and answering (MCQA) [47, 48]. The temperature-weighted softmax scores for each bin give the final action probabilities.
@@
For safety, we restrict the robot and human to work in separate workspaces, such that the human only places objects inside the human workspace and the robot only places block in the robot workspace (i.e., there is no shared workspace). To warm-start the predictor, we allow the human to initially place $3-10$ objects, with eight more to place, for a total of up to $30$ objects per task. We show in Table II that RCIP reduces the plan-wise help rate by $5\%$ and the step-wise help rate by $35\%$ in cooperative manipulation. We use a Franka Emika Panda arm for the robotic manipulation portion of the task. Images of the scene (for both perception and planning) are obtained from an Azure Kinect RGB-D camera.
@@
RCIP $0.85$ $0.85$ 0.95 0.98 0.65
KnowNo [36] $0.85$ $0.85$ $0.96$ 0.98 $0.66$
Simple Set $0.97$ $0.85$ $1.00$ $1.00$ $1.00$
Entropy Set -- $0.66$ $0.46$ $0.45$ $0.06$
No Help -- $0.62$ 00 $0.94$ 00
TABLE II: Results for Cooperative Manipulation. The optimal action miscoverage rate is held fixed between RCIP, KnowNo, and Simple Set for comparing the other metrics.

### v2

@@ IV-E Multi-Step, Multi-Risk Control
$\mathbb{P}^{(\tilde{x},\bar{z})\sim\bar{\mathcal{D}}}(\text{OOD})=\mathbb{P}^{(\tilde{x},\bar{z})\sim\bar{\mathcal{D}}}\Big{(}u^{R*}(z)\notin\mathcal{T}_{\lambda,\theta}(\bar{{x}})\wedge|\mathcal{T}_{\lambda,\theta}(\bar{{x}})|\leq 1\Big{)},$ (24)
@@
In the multi-step, multi-risk setting, we consider a set of sequence-level risk signals $(R_{1},...,R_{k})$ for contexts in $\bar{\mathcal{D}}$ bounded at nominal levels $(\alpha_{1},...,\alpha_{k})$ by all $(\lambda,\theta)\in\hat{\Phi}$ as before. We assume that each risk models an event $E_{k}$, and the loss for each risk $L_{k}$ is an indicator function $\mathbbm{1}[E_{k}]$. We assume that $R_{1}=R_{\text{cov}}$. In addition, since any OOD sequence incurs a task failure, we seek to bound the probability of $E_{k}$ occurring subject to an $R_{1}$ probability of distribution shift (in which case $E_{k}$ can also happen).
@@
Consider a multi-step setting where we use risk calibration with threshold level $(\lambda,\theta)\in\hat{\Phi}$ and the sequence-level score function (21) to generate sequence-level prediction sets and seek help whenever the prediction set is not a singleton. Consider a set of sequence-level risks $(R_{1},...,R_{k})$ bounded at nominal levels $(\alpha_{1},...,\alpha_{k})$, where $R_{1}$ is the miscoverage risk $R_{\text{cov}}$. If the action miscoverage rate is bounded at level $\alpha_{1}$ over the sampling of the calibration set, the new scenarios drawn from $\bar{\mathcal{D}}$ under $\Pi^{R}$ and using the causally reconstructed predicted action set (22) incur at most $\alpha_{1}$ and $\alpha_{k}+\alpha_{1}$ rate of risk for $k\geq 2$ with failure rate $1-\delta$ over the sampling of the calibration set.
@@
$\alpha_{1}+\alpha_{k}$ $\geq R_{1}(\lambda,\theta)+R_{k}(\lambda,\theta)$ (25)
$=\mathbb{E}^{(\tilde{x},\bar{z})\sim\bar{\mathcal{D}}}\Big{[}L_{1}(\tilde{x},\mathcal{T}_{\lambda,\theta}(\tilde{x}),\bar{z})+L_{k}(\tilde{x},\mathcal{T}_{\lambda,\theta}(\tilde{x}),\bar{z})\Big{]}$
@@ V Experiments
Environments. We demonstrate single- and multi-step RCIP in four interactive domains, which exhibit four ways in which a robot planner can be integrated with an intent predictor. First, we consider a multi-hallway setting in which two vehicles must coordinate to reach opposite ends of a room by navigating a set of hallways that are only one vehicle-width wide (see Fig. 3). One vehicle is driven by simulated human that has diverse intent. At the start of the episode, the human vehicle selects one of the hallways at random but does not communicate the hallway to the robot. Next, we investigate human-robot Social Navigation in close-quarters, cluttered household settings in the Habitat 3.0 [47] simulator (see Fig. 4(left)). Third, we show hardware experiments for zero-shot Open-Category Sorting, in which the robot aids the human in sorting common household objects (e.g. books, toys, and fruit) by a mixture of size, shape, and color (see Fig. 1). Finally, we show hardware experiments for a bimanual sorting setting in which the robot must take an object from the human and place it in the human’s intended bin when multiple bins may be appropriate (see Fig. 4(right)). In all settings, since the environment dynamics (1) may evolve at a faster time scale than the human’s intent dynamics (2), the human’s intent is updated once every $T_{z}$ timesteps and is constant otherwise.
@@
Scenario Distribution and Calibration Dataset. RCIP can be used to obtain risk guarantees for an unknown scenario distribution — that is, of environments and human partners — if can can collect i.i.d. samples from it for calibration. We envision that RCIP will enable a robot to interact with an end user (or set of users) through interactive data collection. Then, using the set of FWER-controlling parameters obtained from calibration (cf. section IV), the user may set a level of autonomy for the robot depending on the risk limits of the task. The scenario distribution for each environment is described in the following subsections. Each calibration dataset is generated by random sampling from the environment distribution and from the distribution over human intents. For the simulation environments, we use a pre-trained prediction model using $10k$ random scenarios. For the hallway and Social Navigation environments, the prediction models are trained on a single NVIDIA GeForce RTX 2080 Ti GPU. Pre-training the prediction model takes about $4$ hours per environment. For calibration on hardware, data collection takes about $8$ hours. For all environments, we fix $\delta=0.01$ and use a calibration dataset of size $M=400$. We fix $\alpha_{\text{cov}}=0.15$ for the Hallway and Social Navigation tasks and We fix $\alpha_{\text{cov}}=0.25$ for zero-shot prediction in the Open-Category Sorting and Bimanual Sorting tasks. In all experiments, we evaluate thresholds $\lambda\in[0,1]$ using an evenly spaced grid with $|\hat{\Lambda}|=2000$. We evaluate the model temperature $\theta\in[0.001,10]$ using a log-spaced grid with $|\Theta|=5$. For more details on the scenario distribution, please see Appendix -A.
@@
Baselines. We compare RCIP against similar set-valued prediction approaches. A simple but naive approach for approximated $1-\alpha_{\text{cov}}$ coverage of optimal actions is Simple Set, which ranks actions according to a $1-\alpha_{\text{cov}}$ threshold using the predictor’s raw confidence scores. Actions are sorted by greatest to least confidence, and actions are added to the prediction set in order of the sorted action set until the threshold is reached. To measure the effect of overall uncertainty rather than individual scores, we compare against Entropy Set, which includes the highest overall prediction if the entropy of the distribution predicted actions is below a threshold; if not, then all actions are included in the prediction set, and the robot must ask for help. To evaluate the performance of vanilla conformal prediction against the richer hypothesis space of RCIP, we report results for KnowNo [36]. Similar in spirit but different from our work, KnowNo seeks to maximize coverage of optimal actions but without any guarantees on the human help rate, and assumes model parameters are fixed. Instead of maximizing coverage outright, RCIP balances prediction of optimal actions with limits on the human help rate, providing flexible performance guarantees depending on model parameters. Lastly, we consider No Help as an option, where the predicted action set always contains the predictor’s most-confident action, and the human help rate is identically zero.
@@
V-A Simulation: Hallway Navigation
Figure 4: (Left, Center) Multi-step RCIP is applied in Social Navigation. The human’s trajectory is shown in pink, and the robot’s trajectory is shown in blue. The human’s possible goal objects are shown in orange. (Right) Single-step RCIP is applied in Bimanual Sorting. KnowNo, which generates plans in open-ended language, may generate a plan that is technically correct, but ambiguous to execute for a language-conditioned policy (both the blue and white bin have a pot). RCIP instead guarantees that the human’s intent is satisfied via constraint satisfaction with the intent-conditioned planner.
Figure 5: Baseline comparison for RCIP versus other set-valued predictors for all tasks. RCIP consistently requires less help to achieve a specified plan success rate than other baseline methods. RCIP provides a framework for tuning model parameters to achieve risk control, versus other methods that assume that model parameters are held fixed: KnowNo [36], Simple Set, Entropy Set, and No Help.
[…]
RCIP $0.85$ $0.86$ 0.27 0.98 0.27
KnowNo [36] $0.85$ $0.87$ $0.58$ $0.98$ $0.58$
Simple Set 0.94 $0.86$ $0.54$ $0.98$ $0.54$
Entropy Set -- $0.61$ $0.03$ $0.94$ $0.00$
No Help -- $0.59$ 00 $0.94$ 00
TABLE I: Results for Hallway Navigation. The optimal action miscoverage rate is held fixed between RCIP, KnowNo, and Simple Set for comparing the other metrics.
[…]
1 RCIP $0.86$ $0.27$ 0.98 0.27
No Help $0.59$ 00 $0.94$ 00
2 RCIP $0.87$ $0.04$ 0.98 0.04
No Help $0.68$ 00 $0.94$ 00
3 RCIP $1.00$ $0.00$ 1.00 0.00
No Help $1.00$ 00 $1.00$ 00
TABLE II: Help rate results for Hallway Navigation as the prediction model improves. RCIP is able to keep the plan success rate relatively constant even for poorly performing predictors. As the no-help predictions become better, RCIP also asks for less help.
Autonomous navigation around other autonomous decision-making agents, including humans, requires the robot to recognize scenario uncertainty (whether another agent will turn right or left) with task efficiency (energy spent braking or taking detours). While safety can almost always be guaranteed if each vehicle declares their intent at all times, such communication can be costly, especially if human prompting is involved. In this example (Fig. 3), the robot is asked to navigate to the initial condition of the human’s vehicle without colliding. The set of intents is $\mathcal{Z}=\{1,2,3,4,5\}$, where each intent corresponds to one of the five hallways. The confidence scores for each intent are computed by taking the temperature-weighted softmax scores for each hallway. The final action probabilities are computed according to Eqn. (11). The robot interacts with the human over $T=200$ environment time steps and predicts the human’s intent every $T_{z}=20$ time steps. Additional details on training the intent prediction model are deferred to Appendix -B.
To ensure that the robot reaches its goal state in a minimal amount of time, we permit the robot to prompt the human for their chosen hallway if its optimal action set is not a singleton. We jointly learn the robot and human policies using proximal policy optimization (PPO) [41, 48]. The human and robot PPO policies are trained jointly using $256$ environments and take about $4$ hours to train.
@@ V-A Simulation: Hallway Navigation
We present results for Hallway Navigation in Table I at the fixed coverage level $\alpha_{\text{cov}}=0.85$. We additionally study various model calibration levels and their effect on the RCIP help rate. We present results in Table II
@@
RCIP $0.85$ $0.86$ 0.00 0.90 0.00
KnowNo [36] $0.85$ $0.88$ $0.58$ $0.95$ $0.58$
Simple Set 0.83 $0.86$ $0.56$ $0.94$ $0.56$
Entropy Set -- $0.79$ $0.38$ $0.93$ $0.21$
No Help -- $0.61$ 00 $0.81$ 00
@@
TABLE III: Results for Social Navigation. The optimal action miscoverage rate is held fixed between RCIP, KnowNo, and Simple Set for comparing the other metrics.
V-B Simulation: Social Navigation in Habitat
@@
Habitat [47] is a photo-realistic simulator containing a diverse set of scenes, objects, and humans models for human robotics tasks. In this experiment, a Boston Dynamics Spot robot and human are jointly tasked with navigating to a set of goal objects in sequence, to simulate cleaning up a house (i.e., grabbing various items, such as crackers, cans of soup, etc. as shown in Fig. 4). Each scene contains $5-10$ objects of interest from the YCB dataset [49]. Although the human may initially be out of view of the robot, the robot must find the human and maintain a safe distance of one meter at all times. We simulate the human’s decision making by choosing a high-level intent from the set of objects; here, the set of intents is $\mathcal{Z}=[N_{o}]$, where $N_{o}$ is the number of objects in the scene. The confidence scores for each intended object are computed by taking the temperature-weighted softmax scores for each goal object. The final action probabilities are computed according to Eqn. (11). The robot interacts with the human over $T=600$ environment time steps and selects a new goal object every $T_{z}=100$ time steps.
@@
Since the human’s goal object is not observed by the robot, one naive strategy is to navigate to the human first, then follow the human around the house. However, since the scene is cluttered, remaining too close to the human could impede their progress (e.g. getting in the way) or block the robot, resulting in suboptimal, unsafe behavior. By predicting the human’s motion, the robot is able to better accommodate the human’s task while remaining safe (with high probability) with respect to unsafe interactions. We present results for Social Navigation in Table III. We defer additional discussion, such as the effect of temperature scaling, to Appendix -C.
V-C Hardware: Open-Category Sorting
@@
In this example (Fig. 1), each scenario tasks the robot with helping a human to sort a set of objects by inferring the sorting category for each object. Since the human may have a preference for how the robot sorts the objects, the robot’s trajectory is constrained such that the object must be placed in the correct bin. The human’s intent space $\mathcal{Z}=\{1,2,3\}$, where each number corresponds to one of the bins. The robot interacts with the human over $T=6$ environment time steps, and the human selects a new sorting plan every $T_{z}=1$ time step.
@@
To predict the human’s intent, the robot takes in an image observation of the table and has access to a (vision) language model to process the semantic features of the image. We use GPT-4V (gpt-4-turbo-2024-04-09) [4] to process the image and predict (e.g. the carrot in Fig. 1) a confidence score for each bin via multiple-choice question and answering (MCQA) [50, 51]. The temperature-weighted softmax scores for each bin give the final action probabilities.
@@
For safety, we restrict the robot and human to work in separate workspaces, such that the human only places objects inside the human workspace and the robot only places block in the robot workspace (i.e., there is no shared workspace). To warm-start the predictor, we allow the human to initially place $3-10$ objects, with eight more to place, for a total of up to $30$ objects per task. We show in Table IV that RCIP reduces the plan-wise help rate by $11\%$ and the step-wise help rate by $5\%$ in Open-Category Sorting. We use a Franka Emika Panda arm for the robotic manipulation portion of the task. Images of the scene (for both perception and planning) are obtained from an Azure Kinect RGB-D camera. For more hardware implementation details, please see Appendix -D.
@@
RCIP $0.75$ $0.76$ 0.87 0.94 0.93
KnowNo [36] $0.75$ $0.75$ $0.98$ $0.95$ $0.98$
Simple Set 0.59 $0.76$ $1.00$ $0.96$ $1.00$
Entropy Set -- $1.00$ $1.00$ $1.00$ $1.00$
No Help -- $0.58$ 00 $0.88$ 00
TABLE IV: Results for Open-Category Sorting. The optimal action miscoverage rate is held fixed between RCIP, KnowNo, and Simple Set for comparing the other metrics.
V-D Hardware: Bimanual Sorting
In this experiment (Fig. 4), a bimanual setup with two Franka Panda arms is asked to help the human sort a common household item: frying pan, plushie, wooden block, carrot, or grape. We initialize each bin to contain similar objects; for example, bins can contain other cookware, plushies, wooden blocks, vegetables, or fruit. There are five bins total: two bins are on each side of the table, and one bin is in the middle of the table. The workspace of both arms is limited such that each arm cannot reach the two bins at the other end of the table; thus, there is ambiguity arising from which arm the human wants to help sort the item. Additionally, duplicate categories provide intent ambiguity: each of the initializations is selected at random and may appear up to twice in the five bins. The robot must then interpret a human instruction and infer the set of bins that it can place the object into (Bin 1, Bin 1 or Bin 2, Bin 1, Bin 3, or Bin 5, etc). Therefore, the human’s intent space $\mathcal{Z}$ is the power set of possible bins, i.e., $\mathcal{Z}=\mathcal{P}(\{1,2,3,4,5\})$. Hence, there are $|\mathcal{Z}|=31$ possible intents for this task (when excluding the empty set). The robot interacts over a single time step, so $T=T_{z}=1$. We show in Table V that RCIP reduces the plan-wise help rate by $11\%$ and the step-wise help rate by $5\%$.
Method $1-\alpha_{\text{cov}}$ Plan Success $\uparrow$ Execution Success $\uparrow$ Plan Help $\downarrow$
RCIP 0.75 0.75 0.67 0.85
KnowNo [36] 0.75 0.75 0.60 0.93
Simple Set 0.75 0.75 0.67 0.95
Entropy Set -- 0.94 0.83 0.88
No Help -- 0.40 0.27 0
TABLE V: Results for Bimanual Sorting. The optimal action miscoverage rate is held fixed between RCIP, KnowNo, and Simple Set for comparing the other metrics. We use a random set of 30 scenarios from the test set to evaluate the execution success rate for each method.